OpenStudy (anonymous):

I have absolutely no idea how to do this problem!!!! It says; In the state of North Carolina, if you make less than $12,750 annually you will pay a 10% state tax. If you make at least $12,750 but less than $60,000 annually, you will pay a flat fee of $765 and a 7% tax on every dollar made over $12,750. If you make at least $60,000 annually, you will pay a flat fee of $4,072.50 and a 7.75% tax on every dollar made over 60,000. Write a function that models these tax brackets.

jimthompson5910 (jim_thompson5910):

focus on "if you make less than $12,750 annually you will pay a 10% state tax" and try to make an equation from that

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

10% , x<12,750 ??

OpenStudy (anonymous):

Is that even close?

jimthompson5910 (jim_thompson5910):

let's say x is the amount of money you made before taxes

jimthompson5910 (jim_thompson5910):

if x < 12750, then you take 10% of this to get the amount of tax you have to pay 10% of x = 0.10x

jimthompson5910 (jim_thompson5910):

so x = amount you earn before you pay taxes y = amount of tax you pay the equation is y = 0.10x if x < 12750

jimthompson5910 (jim_thompson5910):

notice how I tacked on the " if x < 12750" that equation is only valid if you make less than $12,750. If you make more than that, then you are not in that tax bracket and another equation applies to you.

jimthompson5910 (jim_thompson5910):

making sense so far?

OpenStudy (anonymous):

I think I'm following you.

OpenStudy (anonymous):

So for if you make at least $60,000, the equation would be 4,072.5+.0775x , and x would greater than or equal to $60,000?

jimthompson5910 (jim_thompson5910):

close, but notice "a 7.75% tax on every dollar made over 60,000"

jimthompson5910 (jim_thompson5910):

saying ".0775x" means you're taxing all of the income at 7.75% when that isn't the case you're taxing the dollars that go over $60,000

OpenStudy (anonymous):

So would that make it only greater than $60,000?

jimthompson5910 (jim_thompson5910):

for example, say you earn $62,000 you'd be taxed a flat $4,072.50 plus an additional amount of 7.75% of $2000 (notice how 62,000 - 60,000 = 2,000)

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

Instead of it being greater than or equal to $60,000, it would be only >. I don't think that's right though.

jimthompson5910 (jim_thompson5910):

saying "at least $60,000" means "60,000 or more" so 60,000 is included

OpenStudy (anonymous):

Okay that makes sense. So what am I doing wrong? I still don't understand.

jimthompson5910 (jim_thompson5910):

let's do this piece by piece

jimthompson5910 (jim_thompson5910):

"if you make less than $12,750 annually you will pay a 10% state tax" the equation here is y = 0.10x if x <12750

jimthompson5910 (jim_thompson5910):

"If you make at least $12,750 but less than $60,000 annually, you will pay a flat fee of $765 and a 7% tax on every dollar made over $12,750" the equation is now y = 765 + 0.07(x - 12750) if 12750 <= x < 60000

jimthompson5910 (jim_thompson5910):

notice how I have x - 12750 this is the amount of dollars you have over 12750 which is taxed at 7%

OpenStudy (anonymous):

Okay, I understand now!

jimthompson5910 (jim_thompson5910):

and finally, " If you make at least $60,000 annually, you will pay a flat fee of $4,072.50 and a 7.75% tax on every dollar made over 60,000." we have this equation y = 4072.50 + 0.0775(x - 60000) if x >= 60000

OpenStudy (anonymous):

WOOHOO!!!! :D I get it now!! Thank you so much!

jimthompson5910 (jim_thompson5910):

we have these 3 equations (with conditions attached to each equation) y = 0.10x if x <12750 y = 765 + 0.07(x - 12750) if 12750 <= x < 60000 y = 4072.50 + 0.0775(x - 60000) if x >= 60000 they then combine to form this piecewise function \[\large f(x) = \begin{cases} 0.10x & \text{if } x < 12,750 \\ 765 + 0.07(x - 12750) & \text{if } 12,750 \le x < 60,000 \\ 4072.50 + 0.0775(x - 60000) & \text{if } x \ge 60,000 \\ \end{cases} \]

jimthompson5910 (jim_thompson5910):

in reality there are a lot more tax brackets, but this gives a somewhat decent idea of how taxes work